# How LLMs Learn to Reason: A Complex Network Perspective

### v2

@@ I.1 Base Model and GRPO Setup
We used the DeepSeek-R1-Distill-Qwen-1.5B model (Guo et al., 2025), trained on the DeepScaleR-Preview-Dataset (Luo et al., 2025a) with their modified version of the veRL open-source engine (Sheng et al., 2024; ByteDance Seed Team and verl community, ). Our Group Relative Policy Optimization (GRPO) hyperparameters followed the first-stage settings, which specified a response length of 8192, as outlined in the DeepScaleR paper (Luo et al., 2025a).
@@ I.2 SFT Dataset for Annealed-RLVR
The SFT dataset for the annealed-RLVR experiment (labeled “Annealed” in Fig. 6) was constructed to focus on difficult problems. First, we trained the base model using GRPO for 100 steps to create a checkpoint (CKPT100), which is in the vicinity of the maximally frustrated state. We then used this checkpoint to generate 8 rollouts for each of the approximately 40,000 problems in the training set. For the roughly 10,000 problems that yielded no correct responses, we generated an additional 52 rollouts, bringing the total to 60 per problem. From this data, we created the SFT dataset by selecting all correct responses for problems where the initial model accuracy was below 10%, which yielded 3,933 trajectories. Finally, we fine-tuned the model on these trajectories for two epochs with a learning rate of $3\times 10^{-5}$. The resulting model served as the new starting point for continuing GRPO training up to step 1,700.
@@ I.3 SFT Dataset for Catastrophic Forgetting
The SFT dataset for the catastrophic forgetting experiment (labeled “SFT+RLVR” in Fig. 4(b)) was designed to specifically target problems forgotten by a more saturated model, using a data generation process that involved multiple checkpoints. The process began with problem filtering: we used a checkpoint from step 600 (CKPT600, deep within the slow-learning stage) to generate 8 rollouts for all training problems. We then identified the subset of problems with no correct responses and generated another 32 rollouts for them, again using CKPT600. Problems that still had no correct responses were selected as the final “forgotten” set. For the second stage of targeted data generation, each problem in this forgotten set was used to generate a total of 96 rollouts: 32 from the base model, 32 from CKPT100, and 32 from CKPT600. We then collected correct responses from problems that had at least 3 correct answers among these 96 rollouts, resulting in 798 trajectories. This SFT dataset was used to fine-tune the model for two epochs with a learning rate of $5\times 10^{-5}$ before resuming GRPO training.
@@ I.4 Computational Cost Analysis
This $O(\eta N)$ SFT cost is significantly smaller than the evaluation cost. Both of these marginal, one-time costs are much smaller than the total iterative cost of the full GRPO training run (e.g., 1,700 steps).
@@
We evaluated checkpoints from our RLVR (“GRPO”) and annealed-RLVR (“Annealed”) training runs on two datasets: 512 randomly selected training problems and the Minerva math dataset (Hendrycks et al., 2021; Lewkowycz et al., 2022). The evaluation parameters were kept consistent with the training configuration (temperature: $0.6$, top-p: $1.0$, response length: $8192$, etc.). From the $512$ rollout results, we calculated the best@k accuracy and the standard error. Fig. 6 illustrates the full performance curves, while Table 1 highlights the results from representative checkpoints.

### v3

@@ I.1 Base Model and GRPO Setup
We used the DeepSeek-R1-Distill-Qwen-1.5B model (Guo et al., 2025), trained on the DeepScaleR-Preview-Dataset (Luo et al., 2025) with their modified version of the veRL open-source engine (Sheng et al., 2025; ByteDance Seed Team and verl community, ). Our Group Relative Policy Optimization (GRPO) hyperparameters followed the first-stage settings, which specified a response length of 8192, as outlined in the DeepScaleR paper (Luo et al., 2025).
@@ I.2 SFT Dataset for Annealed-RLVR
The SFT dataset for the annealed-RLVR experiment (labeled “Annealed” in Fig. 6 of the main text) was constructed to focus on difficult problems. First, we trained the base model using GRPO for 100 steps to create a checkpoint (CKPT100), which is in the vicinity of the maximally frustrated state. We then used this checkpoint to generate 8 rollouts for each of the approximately 40,000 problems in the training set. For the roughly 10,000 problems that yielded no correct responses, we generated an additional 52 rollouts, bringing the total to 60 per problem. From this data, we created the SFT dataset by selecting all correct responses for problems where the initial model accuracy was below 10%, which yielded 3,933 trajectories. Finally, we fine-tuned the model on these trajectories for two epochs with a learning rate of $3\times 10^{-5}$. The resulting model served as the new starting point for continuing GRPO training up to step 1,700.
@@ I.3 SFT Dataset for Catastrophic Forgetting
The SFT dataset for the catastrophic forgetting experiment (labeled “SFT+RLVR” in Fig. 5b of the main text) was designed to specifically target problems forgotten by a more saturated model, using a data generation process that involved multiple checkpoints. The process began with problem filtering: we used a checkpoint from step 600 (CKPT600, deep within the slow-learning stage) to generate 8 rollouts for all training problems. We then identified the subset of problems with no correct responses and generated another 32 rollouts for them, again using CKPT600. Problems that still had no correct responses were selected as the final “forgotten” set. For the second stage of targeted data generation, each problem in this forgotten set was used to generate a total of 96 rollouts: 32 from the base model, 32 from CKPT100, and 32 from CKPT600. We then collected correct responses from problems that had at least 3 correct answers among these 96 rollouts, resulting in 798 trajectories. This SFT dataset was used to fine-tune the model for two epochs with a learning rate of $5\times 10^{-5}$ before resuming GRPO training.
@@ I.4 Computational Cost Analysis
This $O(\eta N)$ SFT cost is significantly smaller than the evaluation cost. Both of these marginal, one-time costs are much smaller than the total iterative cost of the full GRPO training experiment (e.g., 1,700 steps).
@@
We evaluated checkpoints from our RLVR (“GRPO”) and annealed-RLVR (“Annealed”) training experiments on two datasets: 512 randomly selected training problems and the Minerva math dataset (Hendrycks et al., 2021; Lewkowycz et al., 2022). The evaluation parameters were kept consistent with the training configuration (temperature: $0.6$, top-p: $1.0$, response length: $8192$, etc.). From the $512$ rollout results, we calculated the best@k accuracy and the standard error. Fig. 6 of the main text illustrates the full performance curves, while Table 3 highlights the results from representative checkpoints.
@@
Appendix K Supplementary Evidence: Single-Task Dynamics
(a) CoNet
(b) LLM
Figure 17: Single-Task Dynamics We compare the training dynamics of CoNet (a) and the DeepScaleR-1.5B LLM (b) when trained on a single problem. The single-task experiment exhibits: (i) The reward curves (top panels) follow a sigmoidal, phase-transition-like trajectory; (ii) The correct response length (bottom panels) decreases monotonically, converging to the efficient paths. Lines in (a) and (b) are smoothed with a 5-step and 4-step moving average respectively.
We reproduced the single-problem training experiment studied by (Cai et al., 2025) using the DeepScaleR protocol (Luo et al., 2025). Specifically, we selected a single problem from the DeepScaleR-Preview-Dataset (Luo et al., 2025) and repeated the single-task training experiment with their modified version of the veRL (Volcano Engine Reinforcement Learning). As shown in Fig. 17, the single-task dynamics exhibit a sigmoidal learning curve and a monotonically decreasing response length, standing in sharp contrast to the two-stage reward dynamic and V-shaped response length observed during multi-task training. CoNet faithfully reproduces these distinct behaviors when restricted to the same single-task condition.
This single-task training more effectively highlights the phase-transition-like learning mechanism described in the main text. By restricting the system to a single task, we isolate the fundamental phase-transition dynamics of skill acquisition: the model focuses solely on identifying an optimal trajectory among multiple feasible paths. Without the structural requirement to integrate conflicting skills into a shared, sparse backbone, the system avoids the “maximally frustrated state.” Consequently, the response length decreases monotonically as the policy converges to a fixed optimal path, and the reward curve follows a smooth, sigmoidal trajectory characteristic of a continuous phase transition, free from competitive frustration. We also note that the reward curve is influenced by finite-size effects imposed by the maximum response length constraint, this phenomenon remains a key subject for further theoretical investigation.
This capacity to replicate the distinct phenomenologies of both conditions—from the phase transition of isolated learning to the competitive frustration inherent in multi-trajectory integration—provides strong evidence that CoNet serves as a minimal model for investigating the emergent mechanisms of RLVR.